4 | P a g e
Assessment 1.4

CVP analysis in a service business
Edgaria Childcare Ltd plans to open a day care centre at the beginning of next year. A building
has been leased and the company has estimated that the following annual costs will be required:

Salaries — three qualified staff     $128,000
Salaries — two assistants               $64,000
Lease rental (building)                     $62,400
Sundry expenses (all fixed)             $21,600
Supplies                                               780 per child
Meals and snacks                             2080 per child

Based on available space and qualified staff, management believes that the maximum number
of children that can be cared for at the day care centre is 120.

Required

A. Assume that the day care centre can attract 120 children when it opens. Determine the
annual fee per child that must be charged to break even financially.

B. If the company charges $5 980 per child annually, how many children must be enrolled to
break even?

C. Prepare a CVP graph based on a charge of $5 980 per child.

D. If management wants to earn an annual profit of $62 400 from the day care centre and can
charge $5980 per child, how many children must be enrolled?

E. If the day care centre can attract 120 children, how much must be charged per child to earn
a profit of $60 000?

F. If the day care centre can attract 120 children, how much must be charged per child to earn
a profit of $60 000 plus 15% of annual fees?
Answer #1

a.

Salaries 192000
Lease 64000
Sundry Exp 21600
Supplies 780*120 93600
Meals 2080*120 249600
Total 620800
Fees to break even with 120 Students
620800/120 5173.333

b. If $5980 is being charged per year , then total number of students to reach break even shall be = 620800/5980 = 103.8 or 104 students

d. If management wants to earn a profit of $62 400 @5980 per student , then the number of students to be enrolled shall be = (620800+62400)/5989 = 114.07 or 115 students to be enrolled.

e.If day care center wants to earn a profit of $60,000 at 120 students - then the charge per student shall be = (620800 +60,000)/120= $5673.33

f. adding 15% to the fees computed in step e. = $5424.33